OpenStudy (anonymous):

Find the balance on a deposit of $55,000 that earns 2.5% simple interest for 1 year. $1,375 $13,750 $68,750 $56,375

OpenStudy (pixiedust1):

@hartnn

hartnn (hartnn):

simple interest simple formula I = PNR

hartnn (hartnn):

Balance = I + P

OpenStudy (pixiedust1):

It's mah textbook formula!

OpenStudy (anonymous):

wouldn't it be I = PRT

hartnn (hartnn):

your textbook has simple interest formula then not compound interest :)

OpenStudy (pixiedust1):

I see your asking the same type of questions. Is there something in the concept that your not understanding? if so, hartnn may be able to help you. We wanna help you understand the concept, not just individual problems (although we wanna make sure you understand those too)

hartnn (hartnn):

here R = rate = 2.5/100 N = time in years = 1 P = 55000

hartnn (hartnn):

T or N whatever you take as time in years :)

OpenStudy (pixiedust1):

And to be exact, it just says "interest fomula" and goes on to talk about compounding. So I got no idea lol! *Time for a new textbook*

OpenStudy (anonymous):

your looking for I, and once you find I, add it to P